# Generalized Simple Graphical Rules for Assessing Selection Bias

### v3

@@ 4 Estimation and inference by IPW estimators
Based on the identification result in Theorem 2, an unbiased estimator of the ATE is the Hovitz-Thompson (HT) estimator. We also propose an alternative IPW estimator, the Hájek estimator, which is a biased estimator of the ATE but reduces variance compared to the HT estimator [23]. We propose closed-form sandwich estimators of their variances to enable statistical inference of the ATE [24]. We use the two cases in Fig. 1 as demonstrative examples by Corollary 5. We outline the steps of estimation and inference in Appendix C.
@@ 5 Simulation
The true ATE of the treatment is 00, meaning that the medication on average has a zero impact on reducing the risk of malaria.
@@
By Corollary 2, the true ATE is $-0.15$, indicating that the vaccination on average reduces the risk of headache by $15\%$.
@@
The crude selected-sample analysis of Case 2 (red colored in Fig. 4) exhibits biased estimates of the ATE as small as $-0.03$. The $95\%$ confidence intervals contain zero across all three settings. Similar to Case 1, the coverage rate of the true ATE also drops to zero when the sample size increases. The rejection rate is lower than $10\%$ when the sample size is $1000$ or smaller, indicating a type II error rate higher than $90\%$. When the sample size reaches $1000$, the type II error rate is still higher than $50\%$. The explanation is that the decreased chance of reporting whether an individual has a headache among those uninfected partly masks the beneficial indirect effect of the vaccine on reducing headaches through reducing their risk of infection. Therefore, the crude selected-sample analysis fails to reveal the actual beneficial impact of the vaccination on reducing the overall risk of headaches.
Tabel 4 shows the results of the IPW estimators for Case 2 (colored in green and blue). Both methods successfully recover the true ATE around $-0.15$ across all sample sizes. The mean squared error converges to zero. The coverage rates of the true ATE by the $95\%$ confidence interval increase to the intended level. The rejection rates increases to $100\%$ when the sample size is large enough. The power of the two methods with a sample size of $1000$ is greater than that of the crude selected-sample analysis with a sample size of $10000$ by over $10\%$.
@@ 6 Discussion
Our work assesses and extends the utility of existing graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence using d-separation, existing simple graphical rules are inherently coupled with certain identification formulas and applicable to resolving selection bias for a specific subset of selection mechanisms. We show that, some important existing rules [6, 7, 8] can have stringent requirements on covariate adjustment to avoid confounding bias while correcting selection bias. For example, the adjustment for post-treatment variables is generally discouraged by these rules even if it helps separate the selection from the outcome. It is worth noting that while post-treatment variables should not be adjusted for in addressing confounding bias, these variables can indeed be accounted for when dealing with selection bias. We use two concrete examples using causal diagrams to demonstrate the limitations of existing rules and their identification strategies in addressing selection bias.
@@
Specific to the two examples, we theoretically prove the identifiability of the ATE by the g-formula and the IPW method when relevant external information in the full population is available. The g-formula relies on adjusting for the post-treatment upstream cause of selection using a conditional distribution, while the IPW method seeks to balance the distributions of variables by modeling the probability of selection in the full population. The characteristics of these approaches are related to different sufficient conditions that we propose for recovering the ATE. They can be applied to complex observational settings with observed confounders for the relationships between the treatment and post-treatment variables, and unobserved confounders for the relationships between the upstream cause of selection and the outcome. For the IPW method, we also propose two estimators and their variance estimators for the statistical inference of the ATE. We outline the procedures and conducte simulation analysis to verify the performance of our estimators, which correct the serious bias by a complete-case analysis using the selected sample.
@@
Beyond the two examples, our approach to correcting the selection bias validates and suggests possible directions for recovering the ATE when conventional approaches fail. First, researchers have shown that the g-formula and the IPW method are reliable alternatives to the oftentimes biased selected-sample analysis, for addressing selection bias and beyond [25, 14, 17, 18, 19, 11, 20, 21, 22]. Second, it is possible to adjust for post-treatment variables to recover the intended treatment effect without incurring overadjustment bias [26, 27, 28]. The adjustment is in a broader sense that its conditional distribution might be marginalized, or it might be included in a model for the selection, rather than a simple inclusion of it in the outcome regression.
In conclusion, selection bias will continue to be a key topic in epidemiological research. Recent advancements in conceptualizing and addressing selection bias, along with future developments, will bring us closer to a comprehensive understanding of this critical issue.
@@ A.1 The selection-backdoor criterion
The backdoor criterion was proposed to address confounding bias based on DAGs [30]. It was extended to be the selection-backdoor criterion to assess whether conditioning on certain variables $Z$ simultaneously fixes selection bias and confounding bias [6]. The criterion conditions are as follows.
Definition 1 (The selection-backdoor criterion [31]).
@@
(i) Z+ blocks all back door paths from A to Y(ii) A and Z+ block all paths between Z and Y, namely, (ZY|A,Z+)(iii) A and Z block all paths between S and Y, namely, (YS|A,Z)(iv) Z{A,Y}M, and $\begin{split}&\text{(i) $Z^{+}$ blocks all back door paths from $A$ to $Y$}\\ &\text{(ii) $A$ and $Z^{+}$ block all paths between $Z^{-}$ and $Y$, namely, $(Z^{-}\perp\!\!\!\!\perp Y|A,Z^{+})$}\\ &\text{(iii) $A$ and $Z$ block all paths between $S$ and $Y$, namely, $(Y\perp\!\!\!\!\perp S|A,Z)$}\\ &\text{(iv) $Z\cup\{A,Y\}\subseteq M$, and $Z\subseteq T$}.\end{split}$
@@
Following the selection-backdoor criterion based on DAGs, the Generalized Adjustment Criteria were proposed as sufficient and necessary conditions for the assertion of identification or not by certain formulas. Among them, Generalized Adjustment Criterion Type 3 (GACT3) provides the most general conditions for recovering the ATE by a flexible formula that allows the combination of external data and biased selected data [7].
Definition 2 (The Generalized Adjustment Criterion Type 3 (GACT3) [7]).
@@
(a) No element in Z is a descendant in GA¯ of any WA lying on a propercausal path from A to Y(b) All non-causal paths in G from A to Y are blocked by Z and S(c) ZT d-separates Y from S in the proper backdoor graph, where the firstedge of every proper causal path from A to Y is $\begin{split}&\text{(a) No element in $Z$ is a descendant in $G_{\bar{A}}$ of any $W\notin A$ lying on a proper}\\ &\text{causal path from $A$ to $Y$}\\ &\text{(b) All non-causal paths in $G$ from $A$ to $Y$ are blocked by $Z$ and $S$}\\ &\text{(c) $Z^{T}$ d-separates $Y$ from $S$ in the proper backdoor graph, where the first}\\ &\text{edge of every proper causal path from $A$ to $Y$ is removed}.\end{split}$
@@
$\begin{split}\Delta_{\text{ATE},2}=\sum_{z}&(\mathbb{E}[Y|A=1,Z=z,S=1]-\mathbb{E}[Y|A=0,Z=z,S=1])\\ &p(Z\backslash Z^{T}=z\backslash z^{T}|Z^{T}=z^{T},S=1)p(Z^{T}=z^{T}),\end{split}$ (6)
@@ A.3 Simple graphical rules under SWIG by Mathur and Shpitser
Graphical rules based on SWIGs were also proposed for addressing selection bias [8]. It involves two sufficient conditions to analyze whether a conditional average treatment effect (CATE) can be identified by their formula.
Definition 3 (The graphical rules for assessing selection bias [8]).
@@ B.2 Proof of Corollary 1
Let $X$ be an empty set. The conditions (2) hold for the SWIG in Breskin et al.[11] where $Z=W$. By Theorem 1, the g-formula is reduced to
@@ B.6 Proof of Corollary 4
Let $X$ be an empty set. The conditions (4) hold for the SWIG in Breskin et al.[11] where $Z=W$. By Theorem 2, the IPW formula is reduced to
@@
where the second equation is by $S\perp\!\!\!\!\perp A|W$, i.e. there is no direct effect of A on S, for the DAG in Breskin et al.[11]. ∎
@@ Appendix C Steps of IPW estimation and inference
Step 1 Two models for the treatment and selection are required. The propensity score model in the randomized trial, $p(A=1),$ is to our knowledge. Researchers have the flexibility to plug in the true propensity score or an estimated model from observed data. For simplicity of demonstration, we plug in the true propensity score. However, it has been shown that using the estimated propensity score model instead of the true model reduces asymptotic variances of the ATE [23].
@@
Step 3 We calculate the variances for the two IPW estimators by the M-estimation theory [24].
Step 3.1 The first step is to compute a system of equations [32], where the parameters of interest, denoted by $\theta=(\beta_{0},\beta_{1},\mathbb{E}[Y(1)],\mathbb{E}[Y(0)])$, are estimated from
@@
[…]
@@
For both estimators, $\hat{B}(\hat{\theta})$ can be directly computed by plugging in their respective estimating equations (10).

### v4

@@ 4 Estimation and inference by IPW estimators
Based on the identification result in Theorem 2, an unbiased estimator of the ATE is the Hovitz-Thompson (HT) estimator. We also propose an alternative IPW estimator, the Hájek estimator, which is a biased estimator of the ATE but reduces variance compared to the HT estimator [hirano2003efficient]. We propose closed-form sandwich estimators of their variances to enable statistical inference of the ATE [lunceford2004stratification]. We use the two cases in Fig. 1 as demonstrative examples by Corollary 5. We outline the steps of estimation and inference in Appendix C.
@@ 5 Simulation
The true ATE of the treatment is 0, meaning that the medication on average has a zero impact on reducing the risk of malaria.
@@
By Corollary 2, the true ATE is 0.15-0.15, indicating that the vaccination on average reduces the risk of headache by $15\%$.
@@
The crude selected-sample analysis of Case 2 (red colored in Fig. 4) yields biased estimates of the ATE, with point estimates around 0.03-0.03. The $95\%$ confidence intervals contain zero across all three settings. Similar to Case 1, the coverage rate of the true ATE also drops to zero when the sample size increases. The rejection rate is lower than $10\%$ when the sample size is $1000$ or smaller, indicating a type II error rate higher than $90\%$. When the sample size reaches $1000$, the type II error rate is still higher than $50\%$. The explanation is that the decreased chance of reporting whether an individual has a headache among those uninfected partly masks the beneficial indirect effect of the vaccine on reducing headaches through reducing their risk of infection. Therefore, the crude selected-sample analysis fails to reveal the actual beneficial impact of the vaccination on reducing the overall risk of headaches.
Tabel 4 shows the results of the IPW estimators for Case 2 (colored in green and blue). Both methods successfully recover the true ATE around 0.15-0.15 across all sample sizes. The mean squared error converges to zero. The coverage rates of the true ATE by the $95\%$ confidence interval increase to the intended level. The rejection rates increases to $100\%$ when the sample size is large enough. The power of the two methods with a sample size of $1000$ is greater than that of the crude selected-sample analysis with a sample size of $10000$ by over $10\%$.
@@ 6 Discussion
Our work assesses and extends the utility of existing graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence using d-separation, existing simple graphical rules are inherently coupled with specific identification formulas and are therefore applicable only to a subset of selection mechanisms. We show that, some important existing rules [bareinboim2014recovering, correa2018generalized, mathur2023simple] can impose stringent restrictions on covariate adjustment in order to avoid confounding bias while correcting for selection bias. For example, adjustment for post-treatment variables is generally discouraged under these rules, even in settings where such variables help separate selection from the outcome. While post-treatment variables should not be adjusted for when addressing confounding bias, they can nonetheless be appropriately accounted for when dealing with selection bias using methods such as IPW or the g-formula. Using two concrete examples illustrated with causal diagrams, we demonstrate the limitations of existing graphical rules and their associated identification strategies for addressing selection bias.
@@
For the two examples considered, we theoretically establish identifiability of the ATE using both the g-formula and IPW when relevant external information from unselected individuals is available. The g-formula accounts for the post-treatment upstream cause of selection by marginalizing over its conditional distribution, whereas IPW balances covariate distributions by modeling the probability of selection in the full population. These approaches correspond to distinct sufficient conditions that we propose for recovering the ATE. They are applicable in complex observational settings characterized by observed confounding between the treatment and post-treatment variables, as well as unobserved confounding between the upstream cause of selection and the outcome. For the IPW approach, we further propose two estimators along with corresponding variance estimators to enable statistical inference for the ATE. Through simulation studies, we evaluate their performance and show that they substantially reduce the bias introduced by conventional complete-case analyses that rely only on the selected sample.
It should be noted that our simulation studies are intentionally designed to evaluate estimator performance under settings in which the identification assumptions are satisfied, thereby confirming validity when the ATE is identifiable. We acknowledge that the assumption that observed covariates fully d-separate selection from the outcome may be strong and may not hold exactly in practice. Assessing robustness to violations of this assumption is an important direction for future research but is beyond the scope of the present work.
@@
Beyond the two examples presented, our approach to correcting selection bias suggests new directions for recovering the ATE when conventional methods fail. First, prior research has demonstrated that the g-formula and IPW are reliable alternatives to often biased selected-sample analysis, both for addressing selection bias and for broader problems such as generalizability and transportability [robins1986a, imai2010general, cole2010generalizing, stuart2011use, lesko2017generalizing, breskin2018practical, dahabreh2019generalizing, dahabreh2020extending, degtiar2023review]. Second, it is possible to adjust for post-treatment variables in ways that allow recovery of the target treatment effect without inducing overadjustment bias [rosenbaum1984consequences, schisterman2009overadjustment, lu2021revisiting]. Here, “adjustment” is interpreted more broadly: the conditional distribution of a post-treatment variable may be marginalized, or the variable may be incorporated into a model for the selection process, rather than being included directly in the outcome regression.
Importantly, some previously proposed simple graphical rules focus primarily on selected-sample analysis or on estimands other than the ATE, possibly because doing so simplifies covariate selection for adjustment. In practice, however, IPW methods that leverage information beyond the selected sample have long been recognized as effective tools for addressing or mitigating selection bias and recovering the commonly targeted ATE. Selected-sample analysis (i.e., complete-case analysis) alone may be insufficient for recovering the ATE, especially in the presence of type 2 selection bias [lu2022toward].
More broadly, the identification of causal effects in the presence of selection bias depends critically on both the target estimand of interest and the availability of external information. This highlights a key limitation of current practice: while conventional causal diagrams are useful for visualizing and diagnosing the presence of selection bias, they do not encode which variables (exposure, covariates, or outcome) are missing, nor do they indicate whether the treatment effect of interest is identifiable. Integrating causal diagrams for selection bias with graphical models for missing data (often referred to as m-graphs) [mohan2021missing, nabi2020fulllaw] offers a promising direction for overcoming these limitations.
In conclusion, selection bias remains a central challenge in epidemiologic research. Continued progress in its conceptualization, together with the development of robust identification strategies, will bring us closer to a fuller understanding of this fundamental issue.
@@ A.1 The selection-backdoor criterion
The backdoor criterion was proposed to address confounding bias based on DAGs [pearl2009causality]. It was extended to be the selection-backdoor criterion to assess whether conditioning on certain variables $Z$ simultaneously fixes selection bias and confounding bias [bareinboim2014recovering]. The criterion conditions are as follows.
Definition 1 (The selection-backdoor criterion [pearl2014confounding]).
@@
[…]
@@
Following the selection-backdoor criterion based on DAGs, the Generalized Adjustment Criteria were proposed as sufficient and necessary conditions for the assertion of identification or not by certain formulas. Among them, Generalized Adjustment Criterion Type 3 (GACT3) provides the most general conditions for recovering the ATE by a flexible formula that allows the combination of external data and biased selected data [correa2018generalized].
Definition 2 (The Generalized Adjustment Criterion Type 3 (GACT3) [correa2018generalized]).
@@
[…]
@@
[…]
@@ A.3 Simple graphical rules under SWIG by Mathur and Shpitser
Graphical rules based on SWIGs were also proposed for addressing selection bias [mathur2023simple]. It involves two sufficient conditions to analyze whether a conditional average treatment effect (CATE) can be identified by their formula.
Definition 3 (The graphical rules for assessing selection bias [mathur2023simple]).
@@ B.2 Proof of Corollary 1
Let $X$ be an empty set. The conditions (2) hold for the SWIG in Breskin et al.[breskin2018practical] where $Z=W$. By Theorem 1, the g-formula is reduced to
@@ B.6 Proof of Corollary 4
Let $X$ be an empty set. The conditions (4) hold for the SWIG in Breskin et al.[breskin2018practical] where $Z=W$. By Theorem 2, the IPW formula is reduced to
@@
where the second equation is by $S\perp\!\!\!\!\perp A|W$, i.e. there is no direct effect of A on S, for the DAG in Breskin et al.[breskin2018practical]. ∎
@@ Appendix C Steps of IPW estimation and inference
Step 1 Two models for the treatment and selection are required. The propensity score model in the randomized trial, $p(A=1),$ is to our knowledge. Researchers have the flexibility to plug in the true propensity score or an estimated model from observed data. For simplicity of demonstration, we plug in the true propensity score. However, it has been shown that using the estimated propensity score model instead of the true model reduces asymptotic variances of the ATE [hirano2003efficient].
@@
Step 3 We calculate the variances for the two IPW estimators by the M-estimation theory [lunceford2004stratification].
Step 3.1 The first step is to compute a system of equations [reifeis2022variance], where the parameters of interest, denoted by $\theta=(\beta_{0},\beta_{1},\mathbb{E}[Y(1)],\mathbb{E}[Y(0)])$, are estimated from
@@
Step 3.2 The second step is to estimate the asymptotic variance of the parameter estimates. As $n$ goes to infinity, the estimated parameters from equations (10) converge in distribution to a normal distribution, $\hat{\theta}\xrightarrow{d}\mathcal{N}(\theta,V(\theta)/n)$. The covariance matrix, $V(\theta)$, is estimated by a sandwich estimator
@@
[…]